ma scored 273 points last season for her high school basketball team. her points results from a combination of two-points shots and three-points shots. she made a total of 119 shots. how many of each type of shot did ma make?

define 2 variables and write 2 equations and solve them algbraically.

Answer by mathslover(157)

Suppose ma scores x number of 2 point shots and y number of 3 point shots
her total shots = x + y
=> x+y =119......(1)
points scored from her 2 point shots = 2*x
points scored from her 3 point shots = 3* y
total points scored = 2*x + 3*y
=> 2x + 3y = 273 ....(2)
Multipying (1) by 2 on both sides
2x+ 2y = 238 ....(3)
subtract (3) from (2)
y= 35
therefore x = 119 - 35=84
number of 2 point shots 84
number of 3 point shots 35